# What Factors Affect LLMs and RLLMs in Financial Question Answering?

### v3

@@ 4.6 Longer reasoning processes generally lead to better performance
In contrast, for RLLMs, the output of Long CoT indicates that the length of the CoT is not the primary determinant of performance. For instance, DeepSeek-R1-Distill-Qwen-32B generates an excessive number of tokens for easy questions, yet this does not lead to significant performance enhancements compared with Qwen-3-14B. In easy questions, DeepSeek-R1-Distill-Qwen-32B achieves 53.41%, Qwen-3-14B achieves 56.76% in Acc. This phenomenon may arise from the overthinking in RLLMs (Chen et al., 2024), which fails to yield performance gains while substantially increasing computational costs. Moving forward, dynamically adjusting CoT length according to the complexity of the input problem will represent a critical area of research.
@@ 4.7 Performance improves as the parameters of LLMs increase
We further analyze the performance and token consumption of Qwen-3 across various scales and the thinking mode, with the final results illustrated in Figure 3. The data reveal that the performance of LLMs adheres to the scaling law (Kaplan et al., 2020), larger models exhibit enhanced performance and longer output lengths. This suggests that more powerful LLMs engage in more extensive reasoning processes, resulting in superior thinking chain capabilities.
@@
4.8 It’s important to select the optimal alignment language in CLP
We explore the performance differences across various target languages within the CLP method, with the final results presented in Table 4. The findings indicate that different target languages yield varying performance gains, and directly aligning local languages to English through CLP is often not the most effective strategy. Utilizing various language alignments has demonstrated differing impacts on performance (Wang et al., 2025b), likely attributable to LLMs’ comprehension of distinct language families.
Furthermore, different models utilize distinct optimal target languages identified by the CLP method, which may be attributed to varying language capabilities among the models. Although some researchers propose CLSP (Qin et al., 2023) and AutoCAP (Zhang et al., 2024a). CLSP investigates the integration of CLP and self-consistency methods to bridge the gap between different languages. AutoCAP automatically selects languages and automatic weight allocation to obtain the best language combination. This approach incurs substantial operational costs. Therefore, the selection of the best target language based on the multilingual capabilities of each model requires further exploration.
@@ 4.9 Understanding the Advantages of Long CoT
5 Discussion
Furthermore, the Long CoT approach constrains the potential for enhancing performance of RLLMs through agentic frameworks. The advent of agentic reinforcement learning (Agentic RL) enables LLMs and RLLMs to function not only as standalone foundation models but also to be incorporated into complex reasoning and decision-making processes (Luo et al., 2025b; Zhang et al., 2025). Developing a suitable agentic workflow and further training LLMs and RLLMs through agentic reinforcement learning represents an intriguing avenue for improving model performance.
@@ 6 Related Work
With the development of LLMs, more and more researchers focus on the application of LLM in various fields. Improving the performance of LLMs in downstream tasks has consistently been a focal point of researchers’ efforts. Wei et al. (2022) explore the impact of Chain-of-Thought on LLMs and find that forcing thinking of LLMs can significantly improve performance. Inspired by this, researchers further explored the improvement of LLMs. Yao et al. (2023a) propose Tree-of-Thought (ToT), which enhances LLM by considering multiple reasoning paths and self-evaluating. Yin et al. (2023) introduce Exchange-of-Thought (EoT) to enable cross-model communication. In EoT, LLMs could enhance their performance through different network topologies. Zhang et al. (2024b) employs Multi-Perspective Verification and Wrong Information Utilization to prevent LLMs from repeating the same mistakes, thereby significantly enhancing their performance.
Some researchers have introduced complex agentic frameworks to enhance performance by standardizing the processes of LLMs. Yao et al. (2023b) present a novel agent framework ReAct. ReAct can track the interactions of LLMs with the environment and make decisions about the next steps. Li et al. (2023) propose a novel communicative agent framework that enables LLMs to understand tasks through role-playing. Hong et al. (2024) introduce Standardized Operating Procedures (SOPs) into frameworks and propose MetaGPT, which enhances the ability of software engineering in LLMs. Wang et al. (2024) propose S3 Agent to improve the performance of LLMs on multi-modal problems by introducing multiple perspectives information.
@@ 7 Conclusion
The Use of Large Language Models (LLMs)

### v4

@@ 4.6 Longer reasoning processes generally lead to better performance
In contrast, for RLLMs, the output of Long CoT indicates that the length of the CoT is not the primary determinant of performance. For instance, DeepSeek-R1-Distill-Qwen-32B generates an excessive number of tokens for easy questions, yet this does not lead to significant performance enhancements compared with Qwen-3-14B. In easy questions, DeepSeek-R1-Distill-Qwen-32B achieves 53.41%, Qwen-3-14B achieves 56.76% in Acc. This phenomenon may arise from the overthinking in RLLMs (chen2024not), which fails to yield performance gains while substantially increasing computational costs. Moving forward, dynamically adjusting CoT length according to the complexity of the input problem will represent a critical area of research.
@@ 4.7 Performance improves as the parameters of LLMs increase
We further analyze the performance and token consumption of Qwen-3 across various scales and the thinking mode, with the final results illustrated in Figure 3. The data reveal that the performance of LLMs adheres to the scaling law (kaplan2020scaling), larger models exhibit enhanced performance and longer output lengths. This suggests that more powerful LLMs engage in more extensive reasoning processes, resulting in superior thinking chain capabilities.
@@ 4.8 Understanding the Advantages of Long CoT
4.9 It’s important to select the optimal alignment language in CLP
We explore the performance differences across various target languages within the CLP method, with the final results presented in Table 4. The findings indicate that different target languages yield varying performance gains, and directly aligning local languages to English through CLP is often not the most effective strategy. Utilizing various language alignments has demonstrated differing impacts on performance (wang-etal-2025-x), likely attributable to LLMs’ comprehension of distinct language families.
Furthermore, different models utilize distinct optimal target languages identified by the CLP method, which may be attributed to varying language capabilities among the models. Although some researchers propose CLSP (qin-etal-2023-cross) and AutoCAP (zhang2024autocap). These approaches incur substantial operational costs. Therefore, the selection of the best target language based on the multilingual capabilities of each model requires further exploration.
@@ 5 Related Work
With the development of LLMs, more and more researchers focus on the application of LLM in various fields. Improving the performance of LLMs in downstream tasks has consistently been a focal point of researchers’ efforts. wei2022chain explore the impact of Chain-of-Thought on LLMs and find that forcing thinking of LLMs can significantly improve performance. Inspired by this, researchers further explored the improvement of LLMs. yao2023tree propose Tree-of-Thought (ToT), which enhances LLM by considering multiple reasoning paths and self-evaluating. yin-etal-2023-exchange introduce Exchange-of-Thought (EoT) to enable cross-model communication. In EoT, LLMs could enhance their performance through different network topologies. zhang-etal-2024-wrong employs Multi-Perspective Verification and Wrong Information Utilization to prevent LLMs from repeating the same mistakes, thereby significantly enhancing their performance.
Some researchers have introduced complex agentic frameworks to enhance performance by standardizing the processes of LLMs. yao2023react present a novel agent framework ReAct. ReAct can track the interactions of LLMs with the environment and make decisions about the next steps. li2023camel propose a novel communicative agent framework that enables LLMs to understand tasks through role-playing. hong2024metagpt introduce Standardized Operating Procedures (SOPs) into frameworks and propose MetaGPT, which enhances the ability of software engineering in LLMs. 10.1145/3690642 propose S3 Agent to improve the performance of LLMs on multi-modal problems by introducing multiple perspectives information.
@@ 6 Conclusion
Limitations
Due to computational cost constraints, we primarily selected representative methods for our experiments. These methods may not fully reflect the performance of the majority of techniques in financial question answering. To prevent data leakage, we utilized the latest benchmark, FAMMA, for our experiments. However, we cannot guarantee that similar data exists for some questions in the internet data. The challenge of LLMs fitting data requires further exploration in future work.
Acknowledgments
This work was supported by Fund for the Development of Science and Technology (FDCT) of Macau under Grant 0095/2023/RIA2.
Appendix A The Use of Large Language Models (LLMs)
@@
Appendix B Discussion
Furthermore, the Long CoT approach constrains the potential for enhancing performance of RLLMs through agentic frameworks. The advent of agentic reinforcement learning (Agentic RL) enables LLMs and RLLMs to function not only as standalone foundation models but also to be incorporated into complex reasoning and decision-making processes (luo2025agent; zhang2025landscape). Developing a suitable agentic workflow and further training LLMs and RLLMs through agentic reinforcement learning represents an intriguing avenue for improving model performance.
Model Overall (non-thinking) Overall (thinking) Qwen-3-4B 20.35 40.56 Qwen-3-8B 27.66 45.70 Qwen-3-14B 38.09 56.76
Table 5: The performance of the Direct method by Qwen-3-4B/8B/14B in non-thinking and thinking mode.
Appendix C Performance scaling in Qwen-3 non-thinking/thinking mode
We evaluate the performance of Qwen-3 in Direct method under non-thinking/thinking mode, and the final comparison table is shown in Table 5. It can be seen that even small-scale models can benefit from Long CoT.
Method Arithmetic Non-Arithmetic Overall Llama-3.1-8B-Instruct Direct 2.22 12.06 7.76 Zero-shot CoT 6.66 6.89 6.79 Plan-and-Solve 6.66 6.89 6.79 Self-Refine 8.88 6.89 7.76 S3 Agent 2.22 8.62 5.82 GPT-4o-mini Direct 11.11 10.34 10.67 Zero-shot CoT 13.33 17.24 15.53 Plan-and-Solve 17.77 12.06 14.56 Self-Refine 20.00 15.51 17.47 S3 Agent 15.55 3.44 8.73 DeepSeek-R1-Distill-Qwen-32B Direct 26.66 32.75 30.09 Zero-shot CoT 20.00 37.93 30.09 Plan-and-Solve 20.00 32.75 27.18 Self-Refine 17.77 32.75 26.21 S3 Agent 24.44 41.37 33.98 Qwen-3-14B Direct 22.22 37.93 31.06 Zero-shot CoT 33.33 41.37 37.86 Plan-and-Solve 24.44 46.55 36.89 Self-Refine 24.44 36.20 31.06 S3 Agent 22.22 34.48 29.12
Table 6: The results of prompting and agentic methods in FAMMA-LivePro. Bold number presents the best result for these methods on the current model. Light yellow presents the current model is LLM, Light green presents the current model is RLLM.
Appendix D The additional experiment in FAMMA-LivePro
We also conduct a small-scale experiment on FAMMA-LivePro, an additional benchmark independent of the main experiment in this paper, which consists of 103 original questions written by senior financial practitioners. The final experimental results are shown in Table 6. The final results are broadly similar to the conclusions in our paper, so we can draw the same conclusions as the original paper.
Direct Zero-shot CoT Plan-and-Solve Meta-Llama-3.1-8B-Instruct Cost (OpenRouter, Output \$0.05/M) 0.045 0.095 0.110 Overall Acc. 16.50 21.49 22.21 GPT-4o-mini Cost (OpenAI, output \$0.60/M) 0.291 0.502 0.621 Overall Acc. 16.96 39.07 40.46 Qwen-2.5-32B Cost (DeepInfra, output \$0.39/M) 0.280 0.357 0.362 Overall Acc. 44.88 46.11 44.06 DeepSeek-V3 Cost (DeepInfra, output \$0.89/M) 0.737 0.780 0.854 Overall Acc. 58.56 58.71 58.81 DeepSeek-R1-Distill-Qwen-32B Cost (OpenRouter, output \$0.29/M) 1.113 1.132 1.239 Overall Acc. 53.41 53.62 53.26 Qwen-3-14B Cost (OpenRouter, output \$0.24/M) 0.974 0.990 1.032 Overall Acc. 56.76 57.27 56.24
Table 7: The budget results of different methods. Light yellow presents the current model is LLM, Light green presents the current model is RLLM.
Appendix E RLLMs exhibit cost stability across prompting strategies
We analyze the cost of some tokens, and the final results are shown in Table 7. The price of the closed-source model follows the official price, while the price of the open-source model is based on OpenRouter and DeepInfra.
We observe a striking divergence in the cost dynamics between LLMs and RLLMs when subjected to different prompting methods. Specifically, the inference cost of LLMs exhibits high volatility depending on the applied method, whereas RLLMs maintain cost stability. For instance, increasing for Meta-Llama-3.1-8B-Instruct and GPT-4o-mini alongside corresponding accuracy improvements. Conversely, RLLMs display minimal cost variations across prompting methods.